OpenStudy (anonymous):

what is radical 32/4 in simplest radical form?

OpenStudy (mertsj):

\[\sqrt{\frac{32}{4}}=\sqrt{8}=2\sqrt{2}\]

OpenStudy (anonymous):

that isn't a choice. the choices are radical 2 4 radical 2 radical 8 and radical 8/2

OpenStudy (mertsj):

Well, then, choose radical 8. Did you see what I posted?

OpenStudy (precal):

radical 8 is your choice, Mertsj listed in in the step before the last one

OpenStudy (anonymous):

yeah sorry

OpenStudy (precal):

that's ok :)